Suppose that IQ scores have a bell-shaped distribution with a mean of 101 and a standard deviation of 12. Using the empirical rule, what percentage of IQ scores are no more than125?
Answer #1


Solution :

Given that,

Using Empirical rule,

P( - 2< X < + 2) = 95%

P(101 - 2 * 12 < X < 101 + 2 * 12) = 95%

P(77 < X < 125) = 95%

P(X 125) = 1 - 0.025 = 0.975 = 97.5%